1.You have a cylinder of argon gas at 19.8 atm of pressure at 19 oC. The volume of the cylinder is 50.0 L. What would be the volume of the gas if you were to decrease the pressure to 0.974 atm? Assume constant temperature.
2. Helium gas at 22 oC and 1.00 atm occupies a vessel has a volume of 2.54 L. What would be the volume if the gases temperature was decreased to -197 oC? Assume constant pressure
3. A cylinder contains 91.3 g of oxygen gas. If the volume of the cylinder is 8.58 L, what is the pressure of the gas in atmospheres if gas temperature is 21 oC? Calculate the pressure using the ideal gas equation and the van der Waals equation.
4. Calculate the density of H2S gas at 56 oC and 967 mmHg. Obtain the density in grams per liter.
1
Expert's answer
2019-12-12T05:46:22-0500

nRT=PV (Mendeleev-Klapeyrones"s law : ideal gas equation), n - number of moles, R - gase constant 8.314 J/K*mole = 0.0821 l *atm/mol * K , T - absolutly temperature K,

P - pressure, - V - volume ; D - density

(4) 1 moles H2S weight: 34 g;

T(C)+273,15=T(K)

V=nRT/P

V=1 * 0.0821 (l *atm/mol * K ) * 329,15 * K / (967/760 atm)

V = 21,2385 l =21,24 l

D = Weight / volume

D = 34 g / 21,24 l = 1,600 g/l=1,6 g/l

(2) P = const = 1 atm

T1 =295,13K

V1=2,54 l

T2=76,15K

nRT=PV

nR/P = V / T => n =const, R= const, P = const, V/T = const => V1/T1 =V2/T2

V2=V1*T2/T1

V2=0,65537=0,655 l

(1)

nRT=PV, n - const, R -const, T -const, PV -const => P1V1=P2V2

P1=19,8 atm

V1=50 l

P2=0,974 atm

V2=?

V2=P1V1/P2 => 19,8atm*50 l/0,974atm = 1016,4 l

(3) n(O2)=weight/MolarWeight = 91,3g/32g=2,85 mole, R - 0.0821 l*atm/mole*k

V=8,58 l

P=?

T(K)=294,15

nRT=PV

P=nRT/V

P=8,021 atm

Van-Der-Vaals equation:

(p + n2 *a2/V2)(V-n*b)=nRT

As a first approximation, we take the volume V=V1 obtained from the ideal gas equation

V1=nRT/P

V1=V

V1=8,58L

P1=8,021 atm

Pi=n2 *a2/V21

Pi=0,015 atm

second approximation

V2=nRT/(p+pi)+nb

V2=8,5808+0,09006=8,67L

V2=nRT/P

V2=2,85*0,0821*294,15/P

P=7,94

pi2=n2*a2/V22

pi2=0,00199

third(?)

V3=nRT/(p+pi2) +nb=8,7L

P=7,92


Further deviations in approximations of insignificant